If a,b,c in ap ,prove that ab+ac/BC,bc+ba/ca,ca+bc/ab are in ap

Grade:12th pass

3 Answers

mycroft holmes
7 years ago
a,b,c are in AP \Rightarrow \frac{a}{abc} etc. are in AP i.e. \Rightarrow \frac{1}{bc}, \frac{1}{ca}, \frac{1}{ab} are in AP
 
\Rightarrow \frac{ab+bc+ca}{bc}, \frac{ab+bc+ca}{ca}, \frac{ab+bc+ca}{ab} are in AP
 
\Rightarrow 1+ \frac{ab+ca}{bc}, 1+\frac{ab+bc}{ca},1+ \frac{bc+ca}{ab} are in AP
 
\Rightarrow \frac{ab+ca}{bc}, \frac{ab+bc}{ca}, \frac{bc+ca}{ab} are in AP
jagdish singh singh
7 years ago
Very Nice solution bhatt Sir
mycroft holmes
7 years ago
Thanks